The binding of a hormone to three different protein receptors is measured in separate experiments. The table below shows θ, the fractional saturation of each receptor, for various hormone concentrations.

Draw a binding curve that plots the data points for Protein 3. Label the axes of your graph, including the correct units.

What is the Kd for the interaction of the hormone with Protein 3?

Draw binding curves for Proteins 1 and 2 on the same graph. Clearly label which curve represents which of the three proteins.Hormone concentration (nM) Protein Protein 2 Protein 3 0.048 0.11 0.2 0.5 0.71 0.83 0.93 0.2 0.5 0.29 0.5 0.67 0.89 0.95 0.97 0.99 0.17 0.33 0.5 0.8 0.91 0.95 0.98 10 20 50
Answer #1

0.8 0.8 0.6 0.6 prote in1 prote in3 0.4 prote in2 0.4 0.2 0.2 20 20 substrate concentration(nM) substrate concentration(nM)1.2 y -0.9884x+ 0.9972 R2 0.9994 0.8 0.6 bound 0.4 0.2 0 0.2 0.4 0.6 0.8 Bound

from 2nd curve, slope = -1/kd = -0.9884

kd= 1/0.9884 = 1.012*10^-9 M